# Visual-Inertial SLAM as Simple as A, B, VINS

### v2

@@ IV-C3 Robustness to IMU Saturation
IMU saturation can occur during sharp motions or the sensors hitting into an object. While most VINS systems are not robust to IMU saturation, AB-VINS is. To show this, we ran the same simulation as in the previous section, but added saturated IMU measurements to a random axis for 1% of the simulated gyroscope and accelerometer measurements. We use the full VI-BA with depth consistency cost in this experiment. The simulation is run with 1,000 different random seeds both with our robust method for IMU saturation (inflating the noise) and without. It can be seen in Table IX that our robust noise inflation method is successful nearly all of the time, while the standard method fails almost all of the time.
IV-C4 Robustness to Bad Loops
Bad loop closure detections (e.g., from incorrect place recognitions) are typically catastrophic for SLAM systems. On the other hand, our robust $\chi^{2}$ check for loop measurements makes AB-VINS robust to bad loops that may make it past conventional measures to prevent them. To showcase this, using the M3500 dataset, we simulate bad loop measurements for 10% of the loops by adding large random noise to the yaw and position portions of the measurements (standard normal noise in radians for yaw and larger noise for position with standard deviation ranging from 1 to 50 randomly). The results can be seen in Fig. 26, where it can be observed that our robust $\chi^{2}$ check successfully rejects the majority of bad loops, while not using the check results in a visibly worse trajectory. The memory tree top-down optimization method is used in both cases for this experiment.
V Conclusion
We have presented a new visual-inertial SLAM system called AB-VINS, which utilizes three different deep networks. We showed that it is possible, in AB-VINS, to only estimate $a$ and $b$ for monocular depth maps, as well as some other terms ($\mathbf{c}$ and $\mathbf{d}$) to correct the depth with multi-view information instead of each feature position separately. The memory tree, a novel data structure to speed up pose graph optimization, has been introduced. AB-VINS has been shown to have state-of-the-art robustness and efficiency, while also providing dense depth. AB-VINS is a different kind of VINS system – one that heavily relies on deep learning as well as new and improved hand-crafted techniques, and prioritizes efficiency and robustness over accuracy. With this new approach in the open, we hope that researchers can reshape how they do VINS.
VI Acknowledgement
We would like to thank Mingyang Li for his constant support and discussion throughout the development of AB-VINS. We would also like to thank Yuxiang Peng, Saimouli Katragadda, and Chuchu Chen for their help in collecting the low/no excitation dataset. This work was partially supported by the University of Delaware (UD) College of Engineering, the NSF (IIS-1924897, MRI-2018905, SCH-2014264), and Google ARCore.

### v3

@@ V-C3 Robustness to IMU Saturation
IMU saturation can occur during sharp motions or the sensors impacting on an object. While most VINS algorithms are not robust to IMU saturation, AB-VINS is. To show this, we ran the same simulation as in the previous section, but added saturated IMU measurements to a random axis for 1% of the simulated gyroscope and accelerometer measurements. We use the full VI-BA with depth consistency cost in this experiment. The simulation is run with 1,000 different random seeds both with our robust method for IMU saturation (inflating the noise) and without. It can be seen in Table IX that our robust noise inflation method is successful nearly all of the time, while the standard method fails almost all of the time.
V-C4 Robustness to Bad Loops
Bad loop closure detection (e.g., from incorrect place recognition) can be catastrophic for SLAM. Our robust $\chi^{2}$ check for loop measurements makes AB-VINS robust to incorrect loops. To showcase this, using the M3500 dataset, we simulate bad loop measurements for 10% of the loops by adding large random noise to the yaw and position portions of the measurements (standard normal noise in radians for yaw and larger noise for position with standard deviation ranging from 1 to 50 randomly). The results can be seen in Fig. 26, where it can be observed that our robust $\chi^{2}$ check successfully rejects the majority of bad loops, while not using the check results in a visibly worse trajectory. The memory tree top-down optimization method is used in both cases for this experiment.
VI Conclusions
We have presented a new visual-inertial SLAM system called AB-VINS, which utilizes three different deep networks. We showed that it is possible, in AB-VINS, to only estimate $a$ and $b$ for monocular depth maps, as well as some other terms ($\mathbf{c}$ and $\mathbf{d}$) to correct the depth with multi-view information instead of each feature position separately. The memory tree, a novel data structure to speed up pose graph optimization, has been introduced. AB-VINS has been shown to have state-of-the-art efficiency and robustness while also providing dense depth. AB-VINS is a different kind of VINS, one that heavily relies on deep learning as well as new and improved hand-crafted techniques, and prioritizes efficiency and robustness over accuracy. With this new approach in the open, we hope that researchers can re-think how they perform VINS.
Acknowledgement
We would like to thank Dr. Mingyang Li (Google and now Meta) for his constant support and discussion throughout the development of AB-VINS. We would also like to thank our labmates Yuxiang Peng, Saimouli Katragadda, and Chuchu Chen for their help in collecting the low/no excitation dataset. This work was partially supported by the University of Delaware (UD) College of Engineering, the NSF (IIS-1924897, SCH-2014264), and Google.